If sin(θ)=
sqrt 70
14
and 0°<θ<90°, what is cos(θ)?

Answer by stanbon(75887)

If sin(θ)=
sqrt 70
14
and 0°<θ<90°, what is cos(θ)?
-----
By definition, sin = y/r.
So, y = sqrt(70) and r = 14
--
Then x = sqrt[14^2-(sqrt(70))^2] = sqrt(196-70) = sqrt(126) = sqrt(9*14)
= 3sqrt(14)
----
Ans: cos(t) = x/r = 3sqrt(14)/70
----------------------------------
Cheers,
Stan H.
--------------